# Linear Attention for Joint Power Optimization and User-Centric Clustering in Cell-Free Networks

### v1

@@ Abstract
Optimal access point (AP) clustering and power allocation are critical in user-centric cell-free massive MIMO systems. Existing deep learning models lack flexibility to handle dynamic network configurations. Furthermore, many approaches overlook pilot contamination and suffer from high computational complexity. In this paper, we propose a lightweight transformer model that overcomes these limitations by jointly predicting AP clusters and powers solely from spatial coordinates of user devices and APs. Our model is architecture-agnostic to users load, handles both clustering and power allocation without channel estimation overhead, and eliminates pilot contamination by assigning users to APs within a pilot reuse constraint. We also incorporate a customized linear attention mechanism to capture user-AP interactions efficiently and enable linear scalability with respect to the number of users. Numerical results confirm the model’s effectiveness in maximizing the minimum spectral efficiency and providing near-optimal performance while ensuring adaptability and scalability in dynamic scenarios.
@@
Cell-free massive (mMIMO) has emerged as a promising architecture to overcome the limitations of conventional cellular networks [ngo2017cell, ngo2015cell, nayebi2015cell, bjornson2019making]. Unlike traditional cellular systems, where each user equipment (UE) is associated with a single base station, cell-free mMIMO deploys a large number of distributed APs that coherently and jointly serve all UEs within the coverage area. This high level of cooperation offers uniform service quality, mitigates inter-cell interference, and enhances spectral efficiency (SE), particularly in dense or heterogeneous environments. However, realizing these benefits in practice remains challenging. If every AP serves every UE, the resulting system demands substantial fronthaul signaling, creates unnecessary interference from weak AP-UE links, and leads to inefficient use of power and pilot resources [ngo2024ultradense].
To address these challenges, the user-centric paradigm has been introduced in cell-free mMIMO [demir2021foundations, buzzi2017cell, buzzi2025user, bjornson2019new]. In this approach, each UE is served by a subset of the most relevant APs, selected based on spatial proximity or channel quality. By restricting cooperation to the most beneficial APs, the user-centric strategy reduces interference and enhances scalability by lowering the computational and signaling burden. Nevertheless, the performance of user-centric cell-free mMIMO depends critically on the proper selection of AP clusters and transmission powers, which must adapt dynamically to variations in channel conditions, UE mobility, and network configurations. This makes clustering and power control crucial challenges in realizing the full potential of user-centric cell-free mMIMO.
@@ I-A Related work
Given the importance of clustering and power optimization in user-centric cell-free mMIMO, a wide range of solutions have been proposed in the literature [farooq2020accelerated, demir2021foundations, miretti2022closed, shi2025joint, bjornson2013optimal, ammar2021user, di2024deep, liu2016joint, ammar2021downlink, liu2024joint]. These approaches differ in whether they treat clustering and power allocation separately or jointly, and in the methodologies employed, ranging from classical optimization to modern machine learning.
Classical optimization-based methods have long been applied to power allocation, often through iterative algorithms [farooq2020accelerated, demir2021foundations] or closed-form solutions for specific objectives such as the max-min fairness (MMF) problem [miretti2022closed]. While these approaches can achieve near-optimal performance, they typically require multiple iterations to converge, resulting in high computational complexity and poor scalability in dynamic networks.
In parallel, clustering strategies have been studied extensively. Heuristic approaches [shi2025joint, bjornson2013optimal] provide simple rules for grouping users and APs, while optimization-based formulations [ammar2021user] aim for more principled solutions. More recently, deep learning models [di2024deep] have been introduced to automate clustering decisions. Despite their advantages, these methods often lack adaptability: heuristic rules may oversimplify, optimization approaches remain computationally demanding, and deep learning solutions require retraining whenever the number of UEs or APs changes.
Joint clustering and power allocation has also been investigated [liu2016joint, ammar2021downlink, liu2024joint], with the goal of simultaneously determining the serving AP subset and transmission powers. These methods generally outperform separate treatments, but they still suffer from scalability issues and often neglect critical aspects such as pilot contamination. Overall, whether treated separately or jointly, existing solutions remain limited in their ability to adapt efficiently to dynamic network conditions.
@@ I-B Transformer-based power allocation
Recently, Transformers [vaswani2017attention] have been explored to address flexibility in dynamic wireless networks. In [kocharlakota2024pilot], the authors propose a transformer-based model for only down-link (DL) power allocation to handle varying numbers of UEs via unsupervised learning. However, the proposed method requires post-processing and padding, which introduces excessive padding when user loads vary widely, potentially diluting meaningful information, and increases computational load without contributing useful information [dwarampudi2019effects, alrasheedi2023padding], thereby slowing training and affecting convergence. In [chafaa2025transformer], we addressed the flexibility issue in cell-free mMIMO by leveraging a transformer-based model capable of handling varying network configurations without padding or retraining. Unlike [kocharlakota2024pilot], which employs large-scale fading coefficients as input, our model uses only spatial information (AP/UE coordinates) to jointly predict up-link (UL) and DL powers. This design enables efficient adaptation to changes in the number of UEs and APs, while maintaining near-optimal performance for the MMF problem. Nevertheless, the solution did not address AP clustering or pilot contamination, and its complexity still scaled quadratically with the number of users similarly to [kocharlakota2024pilot]. In summary, the key question motivating this work is: How can we overcome these limitations while preserving the advantages of Transformer-based models in dynamic cell-free mMIMO?
@@ I-C Main contributions
Building on our earlier work in [chafaa2025transformer], this paper addresses the limitations previously identified. We introduce a deep learning model, termed exponential linear unit (ELU)-CosFormer, which takes UE and AP coordinates as input and jointly predicts AP clusters and UL/DL power levels. The model employs a customized linear attention mechanism tailored to the spatial and structural characteristics of the prediction task, ensuring scalability and robustness in dynamic network conditions. The learning framework is trained in a supervised manner to maximize the minimum SE, thereby promoting UE fairness while inherently eliminating pilot contamination without requiring additional processing.
The main contributions of this work can be summarized as follows:
  • Joint clustering and UL/DL power control: The proposed model learns to jointly predict the AP clusters serving each UE together with the corresponding UL/DL power levels. This joint design suppresses interference from irrelevant AP-UE links and improves the utilization of limited resources such as power, signaling, and computation. By relying solely on spatial information, the model minimizes overhead and decouples clustering and power control from data detection, while remaining robust to variations in network load.
  • Customized linear attention via ELU-CosFormer: We modify the attention mechanism of CosFormer [qin2022cosformer], a Transformer variant with linear attention complexity, and benchmark it against other linear Transformer architectures [qin2022cosformer, choromanski2020rethinking, wang2020linformer, beltagy2020longformer]. In the proposed variant, the rectified linear unit (ReLU) activation in the attention kernel is replaced with ELU, providing smoother gradients, enhanced numerical stability, and improved convergence under noisy or heterogeneous input conditions.
  • Pilot contamination elimination: The attention mechanism is embedded within a constrained encoder, which enforces that each AP serves no more UEs than the number of available orthogonal pilots. By integrating this constraint directly into the architecture, the model inherently avoids pilot contamination and mitigates interference, without relying on additional pilot-assignment procedures or coordination overhead.
  • Robustness to input uncertainty: We also examine a key deployment aspect that has been overlooked in prior Transformer-based approaches [chafaa2025transformer, kocharlakota2024pilot]: robustness to input uncertainty. In practical systems, AP/UE coordinates are subject to estimation errors; to model this, we perturb the input positions with Gaussian noise of varying intensity. The results demonstrate that the proposed model maintains stable and reliable performance across a broad range of noise levels, highlighting its robustness to imperfect spatial information and its suitability for real-world deployment.
I-D Reproducible research
The code used for model design and training will be provided upon request and will be made publicly available after the review process is complete.
I-E Paper outline and notation
@@
We denote the sets of real and complex numbers by $\mathbb{R}$ and $\mathbb{C}$. Matrices and vectors are written in boldface upper- and lowercase letters, respectively. The superscript $(\cdot)^{\mathsf{H}}$ denotes the Hermitian operator, and $\odot$ indicates element-wise multiplication. Element indices are written as $a_{i}$ for the $i$th entry of vector $\mathbf{a}$ and $a_{i,j}$ for the $(i,j)$th entry of a matrix. The notation $\mathcal{N}_{\mathbb{C}}(\mathbf{m},\mathbf{C})$ refers to a circularly symmetric complex Gaussian variable with mean $\mathbf{m}$ and covariance matrix $\mathbf{C}$. The norm $\lVert\cdot\rVert$ denotes the $\ell_{2}$ vector norm, and $\mathbb{E}[\cdot]$ the expectation operator. The notation $\mathbbm{1}(\cdot)$ refers to the indicator function.
@@
We consider a user-centric cell-free mMIMO network, shown in Figure 1, with $K$ single-antenna UEs and $L$ APs, each having $N$ antennas. Each UE is served by a subset of APs, and each AP can serve at most $\tau_{p}$ UEs to prevent pilot contamination. The network operates under a standard time division duplexing (TDD) protocol [demir2021foundations], where the $\tau_{c}$ symbols of each coherence block are allocated to UL training ($\tau_{p}$), UL data ($\tau_{u}$), and DL data ($\tau_{d}$), with $\tau_{c}\geq\tau_{p}+\tau_{u}+\tau_{d}$. We adopt a narrowband channel model and assume that channels remain constant over a coherence block. The channel vector between AP $l$ and UE $k$ is denoted by $\mathbf{h}_{lk}$ and modeled as [demir2021foundations]:
@@
¡ Refer to caption
Figure 1: Illustration of a user-centric cell-free mMIMO network. Each UE is served by its own cluster of APs [demir2021foundations].
The central processing unit (CPU) computes the estimate of $\mathbf{h}_{k}$ on the basis of received pilot sequences transmitted during the training phase [demir2021foundations]. The minimum mean square error (MMSE) estimate is $\widehat{\mathbf{h}}_{k}=[\widehat{\mathbf{h}}_{1k}^{\mathsf{T}},\ldots,\widehat{\mathbf{h}}_{Lk}^{\mathsf{T}}]^{\mathsf{T}}$ with [demir2021foundations, Sect. IV]
@@ II-A Uplink and downlink spectral efficiency
An achievable UL SE of UE $k$ is given by the use-and-then-forget bound [demir2021foundations, Sect. V]:
@@
with the effective signal-to-interference-plus-noise ratio (SINR) defined as
@@
Here, $p_{k}^{\text{UL}}$ is the UL transmit power of user $k$, $\mathbf{v}_{k}\in\mathbb{C}^{N}$ is the combining vector, and $\mathbf{D}_{k}={\mathrm{diag}}(\mathbf{D}_{k1},\ldots,\mathbf{D}_{kL})$ is a block-diagonal matrix with $\mathbf{D}_{kl}=\mathbf{I}_{N}$ if AP $l$ serves UE $k$, and zero otherwise. The expectation is taken with respect to all sources of randomness. Although the bound in (3) is valid for any combiner $\mathbf{v}_{k}$, we consider the MMSE, given by [demir2021foundations, Sect. V]:
@@
Similarly, the DL SE of user $k$ is [demir2021foundations, Sect. VI]
@@
with effective SINR
@@
Here, $p_{k}^{\text{DL}}$ is the DL power used by the CPU to serve UE $k$ such that $p_{k}^{\text{DL}}=\sum_{l=1}^{L}{p_{k,l}^{\text{DL}}}$, with $p_{k,l}^{\text{DL}}\in\mathclose{[}0,\overline{P}_{l}^{\text{DL}}\mathclose{]}$ being the AP $l$’s transmit power allocated for user $k$; $\overline{P}_{l}^{\text{DL}}$ is the maximum power per AP (when user $k$ is not served by AP $l$, $p_{k,l}^{\text{DL}}=0$); and $\mathbf{w}_{k}\in\mathbb{C}^{LN}$ is the associated unit-norm precoding vector. The MMSE precoder is used [demir2021foundations], which is given by $\mathbf{w}_{k}=\mathbf{v}_{k}/\|\mathbf{v}_{k}\|$.
@@
To ensure fairness among UEs, we adopt the max-min fairness (MMF) problem, widely used in user-centric cell-free mMIMO systems [demir2021foundations, miretti2022closed, chakraborty2019centralized]. In the UL, the problem takes the form [demir2021foundations, Sect. VII]:
@@
where $\overline{P}_{k}^{\text{UL}}$ is the maximum UL power for UE $k$. Similarly, in the DL we have that [demir2021foundations, Sect. VII]:
$\begin{aligned} \max_{\{p_{k}^{\text{DL}}\geq 0\}}\min_{k}&\ \text{\acs{SE}}_{k}^{\text{DL}}\\ \text{subject to}&\quad\textstyle{\sum_{k=1}^{K}p_{k}^{\text{DL}}\leq\sum_{l=1}^{L}\overline{P}_{l}^{\text{DL}}}\end{aligned}$ (10)
where the constraint ensures that the total DL power allocated by the CPU does not exceed the overall network budget.
Although the optimization problems in (9) and (10) can be solved using closed-form solutions [miretti2022closed], iterative solvers [farooq2020accelerated, demir2021foundations], or conventional deep learning models [kim2023survey, mao2018deep, kocharlakota2024pilot], existing methods suffer from two key limitations:
@@
    computational complexity that grows at least quadratically with $K$;
@@
    lack of flexibility with respect to dynamic network loads.
These two drawbacks become critical in large-scale and time-varying networks. Hence, we propose a lightweight linear Transformer model, based on the CosFormer [qin2022cosformer], that exploits only the spatial positions of UEs and APs to adaptively predict both serving clusters and powers while ensuring fairness and scalability under varying UE loads.

### v2

@@ Abstract
Optimal access point (AP) clustering and power allocation are essential in user-centric cell-free massive MIMO systems. However, existing solutions, whether based on deep learning or traditional optimization, lack flexibility to adapt to dynamic network configurations, neglect pilot contamination effects and incur significant computational complexity. In this paper, we propose a lightweight Transformer model that overcomes these limitations by jointly predicting AP clusters and powers solely from the knowledge of spatial coordinates of user devices and APs. Our model is architecture-agnostic to users load, handles both clustering and power allocation without channel estimation overhead, and eliminates pilot contamination by assigning users to APs within a pilot reuse constraint. We also incorporate a customized linear attention mechanism to capture user-AP interactions efficiently and enable linear scalability with respect to the number of users. Numerical results confirm the model’s effectiveness in maximizing the minimum spectral efficiency and providing near-optimal performance while ensuring adaptability and scalability in dynamic scenarios.
@@
[…]
CF massive (mMIMO) has emerged as a promising architecture to overcome the limitations of conventional cellular networks [28, 27, 26, 7]. Unlike traditional cellular systems, where each user equipment (UE) is associated with a single base station, cell-free (CF) mMIMO deploys a large number of distributed APs that coherently and jointly serve all UEs within the coverage area. This high level of cooperation offers uniform service quality, mitigates inter-cell interference, and enhances spectral efficiency (SE), particularly in dense or heterogeneous environments. However, realizing these benefits in practice remains challenging. If every AP serves every UE, the resulting system demands substantial fronthaul signaling, creates unnecessary interference from weak AP-UE links, and leads to inefficient use of power and pilot resources [29].
To address these challenges, the user-centric (UC) paradigm has been introduced in CF mMIMO [13, 8, 9, 6]. In this approach, each UE is served by a subset of the most relevant APs, selected based on spatial proximity or channel quality. By restricting cooperation to the most beneficial APs, the UC strategy reduces interference and enhances scalability by lowering the computational and signaling burden. Nevertheless, the performance of UC CF mMIMO depends critically on the proper selection of AP clusters and transmission powers, which must adapt dynamically to variations in channel conditions, UE mobility, and network configurations. This makes clustering and power control crucial challenges in realizing the full potential of UC CF mMIMO.
@@ I-A Related work
Given the importance of clustering and power optimization in UC CF mMIMO, a wide range of solutions have been proposed in the literature [17, 13, 24, 34, 5, 3, 14, 21, 2, 22]. These approaches differ in whether they treat clustering and power allocation separately or jointly, and in the methodologies employed, ranging from classical optimization to modern machine learning.
Classical optimization-based methods have long been applied to power allocation, often through iterative algorithms [17, 13] or closed-form solutions for specific objectives such as the max-min fairness (MMF) problem [24]. While these approaches can achieve near-optimal performance, they typically require multiple iterations to converge, resulting in high computational complexity and poor scalability in dynamic networks.
In parallel, clustering strategies have been studied extensively. Heuristic approaches [34, 5] provide simple rules for grouping users and APs, while optimization-based formulations [3] aim for more principled solutions. More recently, deep learning models [14] have been introduced to automate clustering decisions. Despite their advantages, these methods often lack adaptability: heuristic rules may oversimplify, optimization approaches remain computationally demanding, and deep learning solutions require retraining whenever the number of UEs or APs changes.
Joint clustering and power allocation has also been investigated [21, 2, 22], with the goal of simultaneously determining the serving AP subset and transmission powers. These methods generally outperform separate treatments, but they still suffer from scalability issues and often neglect critical aspects such as pilot contamination. Overall, whether treated separately or jointly, existing solutions remain limited in their ability to adapt efficiently to dynamic network conditions.
@@ I-B Transformer-based power allocation
Recently, Transformers [35] have been explored to address flexibility in dynamic wireless networks. In [20], the authors propose a Transformer-based model only for down-link (DL) power allocation to handle varying numbers of UEs via unsupervised learning. However, the proposed method requires post-processing and padding, which introduces excessive padding when UE loads vary widely, potentially diluting meaningful information, and increases computational load without contributing useful information [15, 1], thereby slowing training and affecting convergence. In [10], we addressed the flexibility issue in CF mMIMO by leveraging a Transformer-based model capable of handling varying user densities without padding or retraining. Unlike [20], which employs large-scale fading coefficients as input, our model uses only spatial information (AP/UE coordinates) to jointly predict up-link (UL) and DL powers. This design enables efficient adaptation to changes in the number of UEs and APs, while maintaining near-optimal performance for the MMF problem. Nevertheless, the solution did not address AP clustering or pilot contamination, and its complexity still scaled quadratically with the number of users similarly to [20]. In summary, the key question motivating this work is: How can we overcome these limitations while preserving the advantages of Transformer-based models in dynamic CF mMIMO?
@@ I-C Main contributions
Building on our earlier work in [10], this paper develops a scalable geometry-driven optimization framework for UC CF mMIMO systems. The proposed model learns to approximate channel-aware max-min SE optimization policies while significantly reducing computational and signaling overhead.
The main contributions are summarized as follows:
  • Scalable joint clustering and power control: We propose a single end-to-end learning framework that jointly predicts AP clusters and UL/DL power under a max-min fairness objective. Unlike prior methods that rely on repeated channel information for resource allocation, the proposed approach produces near-optimal decisions using only spatial coordinates as input, reducing overhead and decoupling clustering and power control from channel estimation and data detection. Since UE positions capture key propagation and interference characteristics, the model learns this mapping while naturally adapting to varying UE loads.
  • Linear-complexity attention for large-scale CF networks: The proposed model employs a customized linear attention mechanism whose complexity scales linearly with the number of UEs. This makes the framework suitable for dense UC deployments where quadratic-attention Transformers or optimization methods become computationally prohibitive. In addition, the exponential linear unit (ELU)-based mapping improves numerical conditioning and helps preserving weaker but fairness-relevant AP-UE interactions in large heterogeneous networks.
  • Adaptation to varying UE loads without architectural redesign: The architecture is permutation-invariant and operates with different numbers of UEs without architecture modification or retraining, addressing a key limitation of existing deep learning-based resource allocation schemes.
  • Pilot-aware architectural constraint: A pilot-related clustering constraint is embedded directly within the encoder, ensuring that each AP serves no more UEs than the available orthogonal pilots. This mitigates pilot contamination effects without requiring additional pilot assignment procedures.
  • Robustness to imperfect spatial information: We explicitly evaluate robustness to coordinate estimation errors by injecting Gaussian perturbations into UE positions during training and testing. The results demonstrate stable performance under input uncertainty, enhancing practical deployment feasibility.
I-D Paper outline and notation
@@
We denote the sets of real and complex numbers by $\mathbb{R}$ and $\mathbb{C}$. Matrices and vectors are written in boldface upper- and lowercase letters, respectively. The superscripts $(\cdot)^{\mathsf{T}}$ and $(\cdot)^{\mathsf{H}}$ denote the transpose and the conjugate transpose operations, respectively, and $\odot$ indicates element-wise multiplication. Element indices are written as $a_{i}$ for the $i$th entry of vector $\mathbf{a}$ and $a_{i,j}$ for the $(i,j)$th entry of a matrix. The notation $\mathcal{N}_{C}(\bm{\mu},\mathbf{C})$ refers to a circularly symmetric complex Gaussian variable with mean $\bm{\mu}$ and covariance matrix $\mathbf{C}$. The norm $\lVert\cdot\rVert$ denotes the $\ell_{2}$ vector norm, and $\mathbb{E}[\cdot]$ the expectation operator. The notation $\mathbbm{1}(\cdot)$ refers to the indicator function. The $N\times N$ identity matrix and the all-zero vector with $N$ elements are denoted by $\mathbf{I}_{N}$ and $\mathbf{0}_{N}$, respectively.
@@
We consider a UC CF mMIMO network, like the one shown in Figure 1, with $K$ single-antenna UEs and $L$ APs, each having $N$ antennas (in this example, we use $K=3$, $L=8$, and $N=4$). Each UE is served by a subset of APs, called cluster and indicated by the different colors in Figure 1, and each AP can serve at most $\tau_{p}$ UEs to prevent pilot contamination. The network operates under a standard time division duplexing (TDD) protocol [13], where the $\tau_{c}$ symbols of each coherence block are allocated to UL training ($\tau_{p}$), UL data ($\tau_{u}$), and DL data ($\tau_{d}$), with $\tau_{c}\geq\tau_{p}+\tau_{u}+\tau_{d}$. We adopt a narrowband channel model and assume that channels remain constant over a coherence block. The channel vector between AP $l$ and UE $k$ is denoted by $\mathbf{h}_{lk}$ and modeled as [13]:
@@
Figure 1: Illustration of a UC CF mMIMO network, with $L=8$ APs and $K=3$ UEs. Each UE is served by its own cluster of APs.
The central processing unit (CPU), also shown in Figure 1, computes the estimate of $\mathbf{h}_{k}$ on the basis of received pilot sequences transmitted during the training phase [13]. The minimum mean square error (MMSE) estimate is $\widehat{\mathbf{h}}_{k}=[\widehat{\mathbf{h}}_{1k}^{\mathsf{T}},\ldots,\widehat{\mathbf{h}}_{Lk}^{\mathsf{T}}]^{\mathsf{T}}$ with [13, Sect. IV]
@@ II-A Uplink and downlink spectral efficiency
An achievable UL SE of UE $k$ is given by the use-and-then-forget bound [13, Sect. V]:
@@
with UE $k$’s effective signal-to-interference-plus-noise ratio (SINR) defined as
@@
Here, $p_{k}^{\text{UL}}$ is the UL transmit power of user $k$, $\mathbf{v}_{k}\in\mathbb{C}^{N}$ is the combining vector, and $\mathbf{D}_{k}={\mathrm{diag}}(\mathbf{D}_{k1},\ldots,\mathbf{D}_{kL})$ is a block-diagonal matrix with $\mathbf{D}_{kl}=\mathbf{I}_{N}$ if AP $l$ serves UE $k$, and $\mathrm{diag}(\mathbf{0}_{N})$ otherwise. The expectation is taken with respect to all sources of randomness. Although the bound in (3) is valid for any combiner $\mathbf{v}_{k}$, we consider the MMSE, given by [13, Sect. V]:
@@
Similarly, the DL SE of user $k$ is [13, Sect. VI]
@@
where UE $k$’s effective SINR is
@@
Here, $p_{k}^{\text{DL}}$ is the total DL power allocated to serve UE $k$ such that $p_{k}^{\text{DL}}=\sum_{l=1}^{L}{p_{k,l}^{\text{DL}}}$, with $p_{k,l}^{\text{DL}}\in\mathclose{[}0,\overline{P}_{l}^{\text{DL}}\mathclose{]}$ being AP $l$’s transmit power allocated for user $k$; $\overline{P}_{l}^{\text{DL}}$ is the maximum power per AP (when user $k$ is not served by AP $l$, $p_{k,l}^{\text{DL}}=0$); and $\mathbf{w}_{k}\in\mathbb{C}^{LN}$ is the associated unit-norm precoding vector. The MMSE precoder is used [13], which is given by $\mathbf{w}_{k}=\mathbf{v}_{k}/\|\mathbf{v}_{k}\|$. The channel estimation and precoding/combining models adopted above are used to generate channel-aware training data as explained later in Sect. V-A. The proposed learning framework itself does not require instantaneous channel information at the CPU during inference, as it predicts clustering and power decisions directly from UE/ AP spatial information, although channel variations are implicitly embedded during training (see Sect. V-B for further details).
@@
To ensure fairness among UEs, we adopt the MMF problem, widely used in UC CF mMIMO systems [13, 24, 11]. In the UL, the problem takes the form [13, Sect. VII]:
@@
where $\overline{P}_{k}^{\text{UL}}$ is the maximum UL power for UE $k$. Similarly, in the DL we have that:
$\begin{aligned} \max_{\{p_{k,l}^{\text{DL}}\geq 0\}}\min_{k}&\ \text{\acs{SE}}_{k}^{\text{DL}}\\ \text{subject to}&\quad\sum_{k=1}^{K}p_{k,l}^{\text{DL}}\leq\overline{P}_{l}^{\text{DL}}\ \forall l.\end{aligned}$ (10)
The optimization problems in (9) and (10) can be solved using a closed-form solution [24], iterative solvers (e.g., [17, 13]), or deep learning models (e.g., [19, 23, 20]) following these common steps:
@@
    Estimate the large-scale fading coefficients;
@@
    Assign serving AP clusters to the UEs using a clustering algorithm (e.g., the dynamic cooperation clustering (DCC) scheme in [6]);
  3. 3.
    Estimate the channel vectors $\widehat{\mathbf{h}}_{l,k}$ between APs and UEs;
  4. 4.
    Compute the combining and precoding vectors ($\mathbf{v}_{k}$ and $\mathbf{w}_{k}$);
  5. 5.
    Evaluate the SE using (3) or (7);
  6. 6.
    Compute the UL and DL transmit powers by solving (9) and (10).
Irrespective of the adopted methodology, the solution must be obtained in real time, i.e., fast enough to be deployed before the UEs’ positions, load, and spatial distribution change and the clustering and power allocation problems need to be solved again. This stringent latency requirement limits the flexibility of the system with respect to varying UE loads and results in impractical computational overhead in large and dynamic networks, since a new solution must be computed at every coherence block. In particular, the closed-form solution in [24] provides the optimal solution to (9) and (10) for a given fixed clustering configuration. However, it does not address the joint clustering and power allocation problem, and its computational complexity scales cubically with the number of UEs.
To address these limitations, we propose a learning model based on CosFormer [32]. We show that the geographical locations of UEs and APs provide sufficient information to serve as a proxy for jointly determining the clusters and computing the optimal power allocation. We advocate using the UEs’ positions since they capture the key characteristics of propagation channels and interference patterns in the network. The proposed model learns this mapping while naturally handling varying UE loads. Once trained, the model bypasses the repeated execution of the full optimization pipeline (i.e., the six steps listed above) and produces near-optimal decisions through a single forward pass, enabling scalable and real-time adaptation to varying UE loads while preserving fairness.
Furthermore, this framework enables a separation between resource allocation – potentially performed at the CPU – and data detection, carried out at the APs, thereby further improving system scalability. Finally, the proposed approach can also serve as an effective system design tool, as it relies only on coarse information (i.e., the UE locations).